Reports 1-1 of 1 Clear search Modify search
Optical characterization (Optical characterization)
irace, bersanetti, nardecchia, gouaty - 20:09 Tuesday 14 May 2024 (64257) Print this report
OMC scan in CARM_NULL_1F

Figure 1 shows the scan of the OMC performed from 17h05 utc to 17h30 utc, after waiting for 1 hour in CARM_NULL_1F.

Figure 2 shows a zoom of this scan around a Free Spectral Range.

In this FSR, assuming a calibration factor of about 4800 for B1_PD3, we have: ~48 mW on the first order mode, ~24 mW on the second order mode, ~125 mW on the third order modes, ~170 mW on the fourth order mode, around ~40 mW on each of the 5th and 6th order modes, ... Overall, all high order modes are reduced compared to the measurement performed one week ago ( ).

Focusing on the 2nd order mode which appears 4 times during the scan, its power is ranging between 24 and 40 mW which is quite similar to the range of values found last week.

Images attached to this report
Search Help